|Project name|| Main project page|
Previous entry Next entry
Epic failure number 4
On Tuesday, we obtained all of our primers( Lucy, Ricky, Fred) except Ethel. Then we digested our parts (BBa_J37015 = Lux Superpart, BBa_R0062 = Lux promoter,and BBa_J06504 = Old fluorescent protein) from the mini preps of last week and ran electrophroesis. From the gel, we saw that we had plasmid DNA but the digestion did not work properly so we were unable to continue to PCR. Today we redid the digestions with all four enzyme combinations ( ES, EP, XS, XP) with only two samples from each part and ran through electrophoresis. The gel showed we still have plasmids but weird/no digestion patterns. Our next step is to redo the digestions on Tuesday with Axel and the positive control from another group that can do their experiment correctly.